40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Feb 7 09:00:26 2022, Paco, Update, General, Scheduled power outage recovery 
    Reply  Mon Feb 7 15:32:48 2022, Yehonathan, Update, General, Y End laser controller 
    Reply  Mon Feb 7 16:53:02 2022, Koji, Update, General, Scheduled power outage recovery PXL_20220208_001646282.jpgPXL_20220208_001657871.jpg
       Reply  Wed Feb 9 11:56:24 2022, Anchal, Update, General, Bringing back CDS Screenshot_2022-02-09_12-11-33.png
          Reply  Wed Feb 9 13:55:05 2022, Koji, Update, General, Bringing back CDS 
       Reply  Wed Feb 9 16:43:35 2022, Paco, Update, General, Scheduled power outage recovery - Locking mode cleaner(s) 
          Reply  Thu Feb 10 15:41:00 2022, Anchal, Update, General, Scheduled power outage recovery - Locking mode cleaner(s) 
             Reply  Thu Feb 10 17:57:48 2022, Anchal, Update, General, Scheduled power outage recovery - Locking mode cleaner(s) PXL_20220211_021509819.jpg
                Reply  Thu Feb 10 19:03:23 2022, Koji, Update, General, Scheduled power outage recovery - Locking mode cleaner(s) 
                   Reply  Thu Feb 10 19:46:37 2022, Koji, Update, General, Scheduled power outage recovery - Locking mode cleaner(s) 
                      Reply  Thu Feb 10 21:10:43 2022, Koji, Update, General, Video Mux setting reset Screenshot_2022-02-10_21-11-21.pngScreenshot_2022-02-10_21-11-54.png
             Reply  Fri Feb 11 16:09:11 2022, Anchal, Update, General, Scheduled power outage recovery - Input power increased 
                Reply  Mon Feb 14 18:31:50 2022, Paco, Update, General, Scheduled power outage recovery - IMC recovery progress 
                   Reply  Tue Feb 15 19:32:50 2022, Koji, Update, General, Scheduled power outage recovery - IMC recovery progress PXL_20220216_001731377.jpgScreen_Shot_2022-02-15_at_16.18.16.pngPXL_20220216_001727465.jpgScreen_Shot_2022-02-15_at_16.22.16.pngPXL_20220216_002229572.jpg
                      Reply  Tue Feb 15 19:40:02 2022, Koji, Update, General, IMC locking 
                         Reply  Wed Feb 16 15:19:41 2022, Anchal, Update, General, Reconfigured MC reflection path for low power 
                            Reply  Wed Feb 23 15:08:57 2022, Anchal, Update, General, Removed extra beamsplitter in MC WFS path 
                               Reply  Thu Feb 24 14:32:57 2022, Anchal, Update, General, MC RFPD DCMON channel got stuck to 0 
                                  Reply  Thu Feb 24 19:26:32 2022, Anchal, Update, General, IMC Locking 
                                     Reply  Sun Feb 27 00:37:00 2022, Koji, Update, General, IMC Locking Recovery PXL_20220226_093809056.jpgPXL_20220226_093854857.jpgPXL_20220226_100859871.jpgScreenshot_2022-02-26_01-56-31.pngScreenshot_2022-02-26_01-56-47.png
                                        Reply  Sun Feb 27 01:12:46 2022, Koji, Update, General, IMC manual alignment procedure PXL_20220226_100859871.jpg
    Reply  Fri Feb 11 11:17:00 2022, Anchal, Update, General, Scheduled power outage recovery 
       Reply  Mon Feb 14 21:03:25 2022, Koji, Update, General, Scheduled power outage recovery PXL_20220215_025325118.jpg
Message ID: 16658     Entry time: Thu Feb 10 17:57:48 2022     In reply to: 16657     Reply to this: 16659
Author: Anchal 
Type: Update 
Category: General 
Subject: Scheduled power outage recovery - Locking mode cleaner(s) 

Something is wrong with the Video MUX. The system did not turn back on with full functionality. Even though we see the screens as they were before the power shutdown, we have lost control on switching any of the videos. I went to check the wiki page about Video MUX which told be we should be able to see the configuration screen on this link, but the page wasn't opening. I went and removed the power cable and put it back in. That brought back the configuration page. Still, I could not change any of the video feeds however this time, I could see the EPICS channel values (like C1:VID-QUAD1_4) change. I tired to go to the configuration page and change the matrix values from the control tab there. I found out that the matrix was mislabeled and while making the changes, I started seeing blue screen on QUAD1_3 (where MC2T was set before). I set the QUAD1_3 (output 23) to MC2T (input 16), but no change. The EPICS values are also set properly, so I don't understand the reason behind blue screen. The same happened when I tried to use:

~>/opt/rtcds/caltech/c1/scripts/general/videoscripts videoswitch3 QUAD1_3 MC2T

Weirdly, this caused the QUAD1_4 screen to go blue. Running following had no effect:

~>/opt/rtcds/caltech/c1/scripts/general/videoscripts videoswitch3 QUAD1_4 MCR

So, I'm not sure what to do. This really needs to be fixed! I wanted to see teh MC2F camera so that I can align IMC, that was the whole reason for this rabit hole. Help needed.

Attachment 1: PXL_20220211_021509819.jpg  2.102 MB  | Hide | Hide all
PXL_20220211_021509819.jpg
ELOG V3.1.3-